OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge \frac{d}{dx} \cos \sqrt{\sin (\tan x)}\] sounds complicated...

OpenStudy (lgbasallote):

im gonna attempt this...

OpenStudy (lgbasallote):

\[-\sin \sqrt{\sin (\tan x)}\]

OpenStudy (lgbasallote):

that's the first term

OpenStudy (anonymous):

Yes..

OpenStudy (lgbasallote):

\[\frac{1}{2\sqrt{\sin (\tan x)}} \times \cos (\tan x) \times \sec^2 x\]

OpenStudy (anonymous):

Yes you are right lgba..

OpenStudy (lgbasallote):

\[\frac{\sec x \tan x\sin \sqrt{\sin (\tan x)} }{2\sqrt{\sin (\tan x)}}\] is that right?

OpenStudy (anonymous):

Sec^2 or sec??

OpenStudy (lgbasallote):

iguess this wasnt as bad as i thought :)

OpenStudy (lgbasallote):

i cancelled cos

OpenStudy (lgbasallote):

oops wait

OpenStudy (lgbasallote):

yeah i cant do that

OpenStudy (lgbasallote):

\[\frac{\sec^2 x \cos (\tan x) \sin \sqrt{\sin (\tan x) }}{2\sqrt{\sin (\tan x)}}\]

OpenStudy (lgbasallote):

is that better?

OpenStudy (anonymous):

Yes this is the best..

OpenStudy (lgbasallote):

thanks :D

OpenStudy (anonymous):

\[\large \color{blue} {\textbf{Welcome Lgba} \quad {\huge \color{green}\checkmark}}\]

OpenStudy (anonymous):

Welcome me to Latex dear..

OpenStudy (lgbasallote):

\(\color{#c86d3a}{\text{T}}\)\(\color{#bf1678}{\text{H}}\)\(\color{#36fdbf}{\text{A}}\)\(\color{#89ec23}{\text{N}}\)\(\color{#b85e2b}{\text{K}}\)\(\color{#d223a1}{\text{S}}\)\(\color{#127953}{\text{!}}\)\(\color{#f8db7b}{\text{!}}\)\(\color{#6b9572}{\text{!}}\)\(\color{#be5e3b}{\text{!}}\)\(\color{#1ab17b}{\text{}}\)